### v1

# Optimal RIS Placement in a Multi-User MISO System with User Randomness

Abstract
It is well established that the performance of reconfigurable intelligent surface (RIS)-assisted systems critically depends on the optimal placement of the RIS. Previous works consider either simple coverage maximization or simultaneous optimization of the placement of the RIS along with the beamforming and reflection coefficients, most of which assume that the location of the RIS, base station (BS), and users are known. However, in practice, only the spatial variation of user density and obstacle configuration are likely to be known prior to deployment of the system. Thus, we formulate a non-convex problem that optimizes the position of the RIS over the expected minimum signal-to-interference-plus-noise ratio (SINR) of the system with user randomness, assuming that the system employs joint beamforming after deployment. To solve this problem, we propose a recursive coarse-to-fine methodology that constructs a set of candidate locations for RIS placement based on the obstacle configuration and evaluates them over multiple instantiations from the user distribution. The search is recursively refined within the optimal region identified in each stage to determine the final optimal region for RIS deployment. Numerical results are presented to corroborate our findings.
I Introduction
Reconfigurable intelligent surfaces (RIS) have emerged as a promising solution to enhance the performance of a network[1, 2, 3]. An RIS typically operates as a passive reflector that is connected to a smart controller and is capable of adjusting its reflection coefficients based on channel state information (CSI), if available, so that the desired signals are combined constructively and the interfering signals destructively at the receivers[4]. In many cases, its main role in a network is to mitigate blockages caused by buildings and other obstacles[5, 6], a capability that is particularly valuable for higher-carrier-frequency systems, which are known to suffer from severe blockage issues [7, 8]. Given that improving system performance and coverage is the primary utility of an RIS, finding the right place to deploy it in a network is one of the most important and practical problems to address[9, 10].
I-A Related Work
Recent literature on RIS placement addresses the problem from various perspectives. [11] proposes a joint optimization of RIS placement, base station (BS) beamforming, RIS reflection coefficients, and CSI error to maximize the weighted sum rate (WSR) of the system for fixed user locations. In addition, [12] discusses optimizing placement over the coverage area for the special case of indoor communications at THz frequencies. Furthermore, [13] postulates placement optimization over the probability of coverage for random users, BSs, and obstacles, but produces mostly theoretical results and does not engage in performance maximization. [14, 15, 16] also consider the placement problem for specific applications such as high-speed trains, aerial backhaul systems, and unmanned aerial vehicle (UAV) communications.
The performance of an RIS-assisted network is enhanced by jointly optimizing the BS beamforming and the reflection coefficients at the RIS (called joint beamforming). However, since a single beamformer and phase configuration cannot effectively serve all possible user combinations from a distribution, this optimization is mostly performed after the system is deployed. [17] presents one of the most cited joint beamforming algorithms that maximizes WSR. Some other strategies are listed in [18, 19, 20]. The machine learning based approaches are not considered because of their dependence on vast training data, which we might not be able to provide in practical scenarios.
Despite extensive existing research on RIS-based systems, no existing work optimizes RIS placement by maximizing coverage, while also accounting for performance, under user randomness.
I-B Overview of Methodology and Contributions
To address this practical yet unexplored problem, we propose to maximize the expected minimum signal-to-interference-plus-noise ratio (SINR) of the system with respect to the user distribution, under the assumption that joint beamforming is performed after deployment of RIS at its optimal location. The max-min objective inculcates fairness by extending the coverage of the BS to most of the network. However, it is impossible to solve this optimization directly, since the computation of the objective requires knowledge of the beamformer and reflection coefficients, which in turn can only be practically determined after deployment.
Hence, we construct a discrete set of candidate RIS locations for each instantiation of a set of users from the user distribution. For each of these candidates, joint beamforming is performed to compute the optimal beamformer and phases. Using these, for each instantiation of users, the candidate location that maximizes the minimum SINR of the system is chosen to be a feasible solution. Then, this search is further refined within the optimal region identified in each recursive call to determine the final optimal region for RIS deployment.
Figure 1: RIS-assisted MU-MISO system with obstacles.
The proposed technique is the first work, to the best of our knowledge, that simultaneously achieves multiple typical practical requirements of an RIS-assisted system, such as maximizing performance by exploiting the spatial variation of user density, enhancing coverage by exploiting the obstacle configuration, and computing the optimal RIS location without the deterministic user locations. The superiority of the proposed scheme was validated using simulations.
II System Model
Consider a multi-antenna BS at $\mathbf{o_{b}}\in\mathbb{R}^{2}$ serving multiple cells (assumed circular for simplicity) as illustrated in Fig. 1. We assume that all cells operate using orthogonal frequencies, which makes it sufficient to investigate just one cell.
In the cell under investigation, a multi-user multiple-input single-output (MU-MISO) system is considered where the BS with $M$ antennas serves multiple single-antenna users. We also assume that $I$ obstacles hinder signal delivery from the BS. For modeling purposes, these obstacles are represented by two main types: (1) circular obstacles such as pillars and (2) wall-type obstacles. Let $\{\mathbf{o_{o}}\}\in\mathbb{R}^{2}$ denote the locations of the centers of all obstacles. Circular-type obstacles are associated with another set to denote their radii $\{r_{ci}\}$. In contrast, wall-type obstacles are associated with their lengths $\{l_{wi}\}$ and orientations $\{\theta_{wi}\}$. Furthermore, we assume that all obstacles are identical about the z-axis as shown in Fig. 1 (which makes it sufficient to investigate just with its x, y coordinates). Thus, $(\{\mathbf{o_{o}}\},\{r_{ci}\},\{l_{wi}\},\{\theta_{wi}\})$ together is defined as the “obstacle configuration.” From the practical standpoint, this information is fully available to us once the area of deployment (like a specific school) is decided (then the positions of the buildings, pillars, etc. in the school are known). Thus, in this work, we assume that the obstacle configuration is fully known.
On the other hand, practically, we would not know the exact locations of users ahead of deployment. We might only know the spatial variability in user density across the cell. This is mathematically formulated as the user distribution. Unlike conventional spatial distributions like the 2D Gaussian or uniform models, the Poisson Point Process (PPP) provides greater versatility in representing the spatial variability of users in dense environments [13]. Hence, the positions of the users are represented by a homogeneous PPP $\Psi_{u}=\{\mathbf{o_{u}}\}\in\mathbb{R}^{2}$ with density $\lambda_{u}$. Let $K$ denote the number of users in the system corresponding to a specific instance of this PPP.
The direct link from BS to user $k\in\{1,\ldots,K\}$ is modeled using the Rayleigh fading model as
$\mathbf{h}_{b{u}_{k}}=\sqrt{\beta_{bu_{k}}}\tilde{\mathbf{h}}_{b{u}_{k}},$ (1)
where $\beta_{bu_{k}}$ and $\tilde{\mathbf{h}}_{b{u}}\in\mathbb{C}^{M\times 1}$ denote the path loss and the unit-variance circular symmetric complex Gaussian (CSCG) component of the direct link, respectively.
Consider an RIS with $N$ reflecting elements located at $\mathbf{o_{r}}\in\mathbb{R}^{2}$. The links between the BS and RIS and between RIS and the $k^{th}$ user are modeled using the Rician fading model, which are respectively given as
$\mathbf{H}_{br}$ $=\sqrt{\beta_{br}}\left(\sqrt{\frac{T_{1}}{1+T_{1}}}\mathbf{\hat{H}_{br}}+\sqrt{\frac{1}{1+T_{1}}}\tilde{\mathbf{H}}_{b{r}}\right),$ (2)
$\mathbf{h}_{ru_{k}}=\sqrt{\beta_{ru_{k}}}\left(\sqrt{\frac{T_{2}}{1+T_{2}}}\mathbf{\hat{h}_{ru_{k}}}+\sqrt{\frac{1}{1+T_{2}}}\mathbf{\tilde{h}}_{ru_{k}}\right),$ (3)
where $\beta_{br},\beta_{ru_{k}}$ denote the path loss components, $\tilde{\mathbf{H}}_{b{r}}\in\mathbb{C}^{N\times M}$ and $\mathbf{\tilde{h}}_{ru_{k}}\in\mathbb{C}^{N\times 1}$ the unit-variance CSCG components of the links, $\mathbf{\hat{h}_{ru_{k}}}=\mathbf{a}_{M}(\zeta_{k})$ and $\mathbf{\hat{H}_{br}}=\mathbf{a}_{N}(\vartheta)\mathbf{a}_{M}^{H}(\psi)$ the deterministic components of the links. Here, $\mathbf{a}_{i}$ is the steering vector of size $i$ for $i\in\{N,M\}$ and $\vartheta,\psi,\zeta_{k}$ are the angular parameters. $T_{1}$ and $T_{2}$ are the Rician parameters that govern the balance between the steering vector components and the CSCG components. The vectorized reflection matrix of the RIS is given by
$\boldsymbol{\theta}$ $=[\theta_{1},\ldots,\theta_{N}]^{H},$ (4)
where $\theta_{n}=e^{j\varphi_{n}}$ is the phase-shift introduced at the $n^{th}$ element of RIS, for some $1\leq n\leq N$. Let the signal transmitted from the BS be $\mathbf{x}=\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}$, where $\mathbf{w}_{k}\in\mathbb{C}^{M\times 1}$ is the beamforming vector and $s_{k}$ the unit-power transmit information for the $k^{th}$ user. For computational convenience, let us define $\mathbf{H}_{k}=\text{diag}(\mathbf{h}_{{ru}_{k}}^{\mathrm{H}})\mathbf{H}_{br}$. Then, the received signal at user $k$ can be expressed as
$y_{k}=\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\sum_{k=1}^{K}\mathbf{w}_{k}s_{k}+n_{k},$ (5)
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN at the $k^{th}$ user. The corresponding SINR at user $k$ is
$\gamma_{k}=\frac{\left|\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\mathbf{w}_{k}\right|^{2}}{\sum_{i\neq k}\left|\left(\mathbf{h}_{{bu}_{k}}^{\mathrm{H}}+\boldsymbol{\theta}^{\mathrm{H}}\mathbf{H}_{k}\right)\mathbf{w}_{i}\right|^{2}+\sigma^{2}}.$ (6)
III Problem Formulation
Clearly, for a system as shown in Fig. 1, an RIS is required to extend the coverage of the BS to most of the cell. This must be done, also ensuring good performance. This problem is formulated as
$\mathbf{P1.1}:\quad\max_{\mathbf{o_{r}}}$ $\quad\mathbb{E}_{\mathbf{o_{u}}}\left[\min\gamma_{k}\right]$
in which we assume that joint beamforming is optimized using
$\mathbf{P1.2}:\quad(\boldsymbol{\theta},\mathbf{W})=\operatorname*{arg\,max}$ $\sum_{k=1}^{K}\frac{1}{K}\log(1+\gamma_{k})$
[…]
$|\theta_{n}|=1,\forall n=1,\cdots,N.$ (7b)
where $\mathbf{W}=[\mathbf{w_{1}},\cdots,\mathbf{w_{K}}]\in\mathbb{C}^{M\times K}$ is the accumulated beamformer matrix. The objective in $\mathbf{P1.2}$ is the WSR with equal priority assigned to all users. Eq. (7a) enforces the total transmit power constraint $P_{max}$ for the system. Eq. (7b), on the other hand, defines the feasible set for the RIS reflection coefficients by enforcing unit amplitude at the $N$ elements.
In the proposed problem, we intend to find the optimal location for the RIS that maximizes the expected SINR corresponding to the most affected user (max-min) to enhance the coverage of the system. Let us denote this, $\mathbf{P1.1}$, as the “outer optimization”. It should be noted that the ‘expectation’ is taken over the user distribution. We perform this optimization assuming that joint beamforming is performed after the RIS is deployed at its optimal location. This part of the problem, $\mathbf{P1.2}$, is denoted as the “inner optimization”.
Many researchers consider this multilayered structure of the formulation to be the ‘ideal’ formulation that is most relevant practically. In the next section, we propose an elegant technique to solve this practical problem.
IV Proposed Solution
In this section, we first consider the joint beamforming problem $\mathbf{P1.2}$, then the outer optimization problem $\mathbf{P1.1}$.
IV-A The Inner Optimization Problem
IV-A1 Transformed Problem
Using the Lagrangian dual transform technique for closed-form fractional programming [22], problem $\mathbf{P1.2}$ is reduced to a sum-of-ratios problem by introducing a set of auxiliary variables $\boldsymbol{\alpha}=\{\alpha_{k}\}$. Then, the quadratic transform technique [21] is used to decouple the SINR terms (ratios) in the sum-of-ratios problem by introducing another set of auxiliary variables $\boldsymbol{\beta}=\{\beta_{k}\}$, to transform the problem into
$\mathbf{P2.1}:\quad\max_{\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta}}$ $f(\boldsymbol{\alpha},\boldsymbol{\beta},\mathbf{W},\boldsymbol{\theta})$
$=$ $\sum_{k=1}^{K}\Big[\frac{1}{K}(\log(1+\alpha_{k})-\alpha_{k})$
$+$ $2\sqrt{\frac{1}{K}(1+\alpha_{k})}\mathcal{R}e\Big\{\beta_{k}^{*}(\mathbf{h}_{{bu}_{k}}^{H}+\boldsymbol{\theta}^{H}\mathbf{H}_{k})\mathbf{w}_{k}\Big\}$
- $|\beta_{k}|^{2}\Big(\sum_{i=1}^{K}|(\mathbf{h}_{{bu}_{k}}^{H}+\boldsymbol{\theta}^{H}\mathbf{H}_{k})\mathbf{w}_{i}|^{2}+\sigma^{2}\Big)\Big]$
$s.t.\quad$ $\sum_{k=1}^{K}\|\mathbf{w}_{k}\|_{2}^{2}\leq P_{max},$ (8a)
$|\theta_{n}|=1,\forall n=1,\cdots,N.$ (8b)
IV-A2 Iterative Optimization
In order to solve this transformed problem, assuming full CSI, we adopt an iterative optimization strategy proposed in [17]. Let $t$ denote the current iteration. The auxiliary variables are updated with their optimal values obtained by setting $\frac{\partial f}{\partial\alpha_{k}}=0$ and $\frac{\partial f}{\partial\beta_{k}}=0$. These optimal values for $\alpha_{k}$ and $\beta_{k}$ are
$\alpha_{k}(t)$ $=\frac{(\zeta_{k}(t-1))^{2}+\zeta_{k}(t-1)\sqrt{(\zeta_{k}(t-1))^{2}+4}}{2},$ (9)
$\beta_{k}(t)$ $=\frac{\mu_{k}(t-1)(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t-1)\mathbf{H}_{k})\mathbf{w}_{k}(t-1)}{\sum_{i}|(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t-1)\mathbf{H}_{k})\mathbf{w}_{i}(t-1)|^{2}+\sigma^{2}},$ (10)
where $\zeta_{k}(t)=\sqrt{K}\mathcal{R}e\left\{\beta^{*}_{k}(t)(\mathbf{h}_{bu_{k}}^{H}+\boldsymbol{\theta}^{H}(t)\mathbf{H}_{k})\mathbf{w}_{k}(t)\right\}$ and $\mu_{k}(t)=\sqrt{\frac{1}{K}(1+\alpha_{k}(t))}$. Here, $\mathcal{R}e\{.\}$ represents the real part. The update rule for $\mathbf{W}$ is given as
$\mathbf{w}_{k}(t)$ $=\frac{1}{L(t)-2\lambda(t)}(L(t)\hat{\mathbf{w}}_{k}(t-1)-\mathbf{g}_{k}(t)),$ (11)
[…]
where $\hat{\mathbf{w}}_{k}(t)=\mathbf{w}_{k}(t)+\epsilon(\mathbf{w}_{k}(t)-\mathbf{w}_{k}(t-1))$ is the extrapolated beamformer for some $\epsilon\geq 0$ and $\mathbf{g}_{k}(t)=-\frac{\partial f_{1}}{\partial\mathbf{w_{k}}}$ in which $f_{1}(\mathbf{W})=f(\boldsymbol{\alpha}(t-1),\boldsymbol{\beta}(t-1),\mathbf{W},\boldsymbol{\theta}(t-1))$. Also, $L(t)$ is set to the Lipschitz constant of $\mathbf{g}_{k}(t)$.
Finally, the update rule for phase coefficients $\boldsymbol{\varphi}=[\varphi_{1},\cdots,\varphi_{N}]^{T}$ (recall $\theta_{n}=e^{j\varphi_{n}}$) is derived based on the second-order Taylor series expansion of $f_{2}(\boldsymbol{\varphi})=f(\boldsymbol{\alpha}(t-1),\boldsymbol{\beta}(t-1),\mathbf{W}(t-1),e^{j\boldsymbol{\varphi}})$ to be of the form
$\boldsymbol{\varphi}(t)\leftarrow\boldsymbol{\varphi}(t-1)-\frac{\nabla f_{2}(\boldsymbol{\varphi}(t-1))}{\kappa},$ (13)
where optimal $\kappa$ is obtained using the strategy listed in [17].
IV-B The Outer Optimization Problem
Since closed-form solutions for $\mathbf{w}$ and $\varphi$ as functions of the channel, valid for all possible instantiations of the user distribution is clearly intractable, we approximate the expectation in $\mathbf{P1.1}$ over multiple realizations of the user distribution.
For each instantiation of a set of users, a discrete set of candidate locations for RIS (of cardinality $T$) is formed using Algorithm 1 assuming that Circle $C$ is the entire cell. The main idea behind the algorithm stems from the necessity for the RIS to have a clear Line of Sight from the BS to be able to forward the message to the user. In addition, the RIS is placed at least $d_{FF}=\frac{2D^{2}}{\lambda_{c}}$ far from the BS (Fraunhofer distance of the BS), so that the RIS does not interfere in the near field of the BS [23]. Here, $\lambda_{c}$ denotes the wavelength corresponding to the frequency of operation and $D$ denotes the overall aperture size of the BS.
Algorithm 1 Proposed Construction of the Candidate set $\mathcal{Q}$
Inputs:
  1. (i)
    An instance of users: $\{\mathbf{x_{u}}\}$,
  2. (ii)
    BS location: $\mathbf{o_{b}}$,
  3. (iii)
    Obstacle configuration: $(\{\mathbf{o_{o}}\},\{r_{ci}\},\{l_{wi}\},\{\theta_{wi}\})$,
  4. (iv)
    Circle $C$.
Output: $\mathcal{Q}$
1:Set $\mathcal{Q}=\varnothing$ and $i=1$.
2:while $i\leq T$ do
3:  Generate random $\mathbf{q}_{i}$ in $(i\text{ mod }4)^{th}$ quadrant of $C$
4:  if No obstacle lies on the direct link from $\mathbf{o_{b}}$ to $\mathbf{q}_{i}$ and $\mathbf{q_{i}}$ is at least $d_{FF}$ away then
5:   Update $\mathcal{Q}=\mathcal{Q}\cup\{\mathbf{q}_{i}\}$
6:   Set $i=i+1$
7:  end if
8:end while
A method to construct a set of feasible ‘min-max’ solutions $\mathcal{S}$ is proposed in Algorithm 2 based on Algorithm 1.
Algorithm 2 Proposed Construction of the Solution set $\mathcal{S}$
Inputs:
  1. (i)
    User distribution,
  2. (ii)
    Circle $C$.
Output: $\mathcal{S}$
1:Create multiple instantiations of user distribution and use Algorithm 1 to get corresponding candidate sets in $C$
2:For each candidate in each of these candidate sets corresponding to each instantiation, perform joint beamforming
3:For each instantiation, a feasible solution is obtained using $\mathbf{P1.1}$, forming a solution set $\mathcal{S}$
IV-C Computing the Final Solution
It is important to observe that the solution set is not closed under linear combinations, i.e. $s_{1}\in\mathcal{S}$ and $s_{2}\in\mathcal{S}\nRightarrow(a.s_{1}+b.s_{2})\in\mathcal{S},\forall a,b\in\mathbb{R}$. This rules out the usual tendency to compute the mean of all optimal locations to find the ‘final solution’ (mean of solutions not necessarily a solution itself).
Algorithm 3 is a recursive method proposed to compute the final solution. The algorithm was developed by observing the spatial distribution of the solutions for a wide variety of scenarios and user distributions. ‘Clusters’ are formed among the solutions in $\mathcal{S}$ by creating quantized versions of them with a suitable step size, $d_{start}$. Then the cluster with the highest number of entries is declared to be the ‘optimal cluster’. Then Algorithm 2 is repeated for a new candidate set with candidates only within the optimal cluster. This process of looking deeper into the optimal cluster is repeated up to the required precision $d_{p}$. Hence, the globally optimal region to deploy the RIS is obtained as the square $\mathbf{\hat{x}_{r}}(1)-d_{p}/2,\mathbf{\hat{x}_{r}}(1)+d_{p}/2)\times[\mathbf{\hat{x}_{r}}(2)-d_{p}/2,\mathbf{\hat{x}_{r}}(2)+d_{p}/2)$.
Algorithm 3 Computing the final solution
Inputs:
  1. (i)
    Solution set: $\mathcal{S}$,
  2. (ii)
    Initial step size: $d_{start}$,
  3. (iii)
    Precision threshold: $d_{p}$.
Output: $\mathbf{\hat{x}_{r}}=\textsc{RecursiveAlgo}(\mathcal{S},d_{start})$
1:procedure RecursiveAlgo($\mathcal{S}_{temp},d$)
2:  if $d<d_{p}$ then
3:   return mode of set $\mathcal{C}_{temp}$
4:  else
5:   Define $\mathcal{C}_{temp}=\{\text{round}(c/d)\cdot d\mid c\in\mathcal{S}_{temp}\}$
6:   Construct $\tilde{\mathcal{S}}$ using Algorithm 2 for a circle $C$ centered at the mode of $\mathcal{C}_{temp}$ and fixed radius $r$
7:   return RecursiveAlgo($\tilde{\mathcal{S}},d/2$)
8:  end if
9:end procedure
V Simulation Results
(a) Scenario 1: Solutions - RL 1.
(b) Scenario 1: Frequency - RL 1.
(c) Scenario 1: Frequency - RL 2.
(d) Scenario 1: Final Solution.
(e) Scenario 2: Final Solution.
(f) Scenario 3: Final Solution.
(g) Scenario 4: Final Solution.
(h) Coverage Analysis.
Figure 2: Optimal RIS placement in different scenarios.
The simulation parameters used are laid out below, followed by a detailed review of the results obtained.
V-A Simulation Setup
The path loss parameters $\beta_{br}$, $\beta_{ru_{k}}$, $\beta_{bu_{k}}$ are assumed to obey the 3GPP standard for Urban Micro cell (Table B.1.2.1-1 in [24]),
$\beta_{i}=22\log_{10}(d_{i})+28+20\log_{10}(f_{c})\text{, }i\in\{br,ru_{k}\},$ (14)
$\beta_{bu_{k}}=36.7\log_{10}(d_{bu_{k}})+22.7+26\log_{10}(f_{c}),$ (15)
where $d_{i}$ represents the Euclidean distance traversed by each link in meters (m) and $f_{c}$ is the carrier frequency in GHz. $f_{c}$ is set as $2.4$ GHz. For example, we get a loss of $\beta_{br}=\beta_{ru_{k}}=55.6$ dB and $\beta_{bu_{k}}=69.3$ dB for $d_{br}=d_{ru_{k}}=d_{bu_{k}}=10$ m. Then the Rician factors are set to $T_{1}=T_{2}=10$ dB. The transmit power is set to $P_{max}=0$ dB unless mentioned otherwise. In addition,
$\sigma^{2}=-174+\log_{10}(B)+NF\quad\text{ dBm},$ (16)
where bandwidth is set to $B=10$ MHz (since we consider single-carrier communication) and receiver noise figure $NF=5$ dB [24]. Also, in this work, we assume that when a link hits an obstacle, the link power is reduced to a negligible level due to the severe penetration loss.
The BS is assumed to be located at $\mathbf{o_{b}}=(80$ m,30,30 m)) at a corner of the cell as in Fig. 1 and the cell radius is fixed as $R=20$ m. Furthermore, other system parameters are set to $M=16,N=100,d_{p}=1$ m for all simulations. Finally, coverage is computed over a large grid of points in the cell (but not on top of any obstacle) with a resolution of $0.1$ m.
V-B Results and Discussion
Four different scenarios are considered for simulation as shown in Fig. 2. The first 3 scenarios assume various positions, sizes, and orientations for obstacles that hinder access to a set of users following a homogeneous PPP with a fixed density $\lambda_{u}$ throughout the circular cell. And finally, a more practical scenario with 4 user hotspots is considered.
Fig. 2(a)-(d) present a top-level view of the complete procedure described in Section IV for Scenario 1. Fig. 2(a) shows a scatterplot of the solution set in recursion level (RL) 1, in which the black objects are the obstacles, the green circle represents the user cluster (the whole cell), and the blue dots represent the solutions. Fig. 2(b) represents the frequency heatmap of a quantized version of this solution set with $d_{start}=2$ m. To achieve $d_{p}=1\text{ m}$, another recursive pass of the algorithm is run as shown in Fig. 2(c). The cluster with the highest RL 2 frequency is declared as the optimal region for RIS deployment (represented by a $d_{p}\times d_{p}$ square) in Fig. 2(d). As shown in Fig. 2(h), deploying RIS at this location extends coverage from 65.76% to 99.68% of the cell. The average WSR of deploying RIS at the center of this cluster is captured in Table I. Note that the expected number of users corresponding to this scenario is $E[K]=\lambda_{u}.\text{Area}=11.3097$.
TABLE I: Obtained results
Scenarios $\mathbf{\lambda_{u}}$ Final Solution Average WSR Coverage
(bps/Hz)
1 0.009 (105 m,31 m) 5.5631 99.68%
2 0.007 (105 m,42 m) 5.8089 97.91%
3 0.009 (113 m,31 m) 5.8040 97.21%
4 $\begin{cases}0.4,&\text{Circle }C_{1}\\ 0.04,&\text{Circle }C_{2}\\ 0.05,&\text{Circle }C_{3}\\ 0.035,&\text{Circle }C_{4}\\ 0,&\text{elsewhere}\end{cases}$ (111 m,41 m) 3.5982 99.69%
A major obstacle group is considered at the center of the cell in Scenario 2, as shown in Fig. 2(e). A lower user density (on average, three fewer users) is assumed. The optimal location and average WSR is given in Table I. As shown in Fig. 2(h), this extends coverage from 66.27% to 97.91% of the cell. Scenario 3 as shown in Fig. 2(f) was designed to be difficult to extend coverage to the entire cell with a single RIS. The algorithm still derives a solution that covers regions shadowed by the major clusters and covers a portion of the region shadowed by the single wall obstacle (i.e., extends coverage to 97.21% of the cell), with a similar WSR performance.
Finally, a more practical Scenario 4 is considered as depicted in Fig. 2(g) with location dependent $\lambda_{u}$ as given in Table I. Here, Circles $C_{1},C_{2},C_{3},C_{4}$ are assumed as shown in Fig. 2(h). The same obstacle configuration as in Scenario 3 is assumed. The final solution shifts from (109 m, 45 m) to (114 m, 42 m) to accommodate the significantly higher user density in the region shadowed by the wall obstacle, highlighting the capability of the proposed method to converge to a location that not only maximizes coverage by covering most of the shadowed regions, but also effectively adapts to spatial variations in user density.
Figure 3: WSR-transmit power characteristic for Scenario 2.
The average WSR vs transmit power characteristics for Scenario 2 is shown in Fig. 3. The average WSR of deploying RIS at the optimal location significantly exceeds the ‘No RIS’ simulation. This is the increase achieved solely because of the extension of coverage by the RIS to almost the full cell. Furthermore, when the RIS is randomly positioned within the cell, and for each of these random locations - averaged over multiple realizations of the user distribution, the resulting average performance remains close to that of the ‘No RIS’ scenario. This observation underscores the necessity for the optimal placement of RIS. The range of WSR performance across scenarios aligns with similar works [11, 12], reaffirming the optimality of the solution obtained in our work. It is also important to note that this is achieved despite having lower $M$ ($16<32$), practical number of users (average $K\approx 11\gg 2$), different obstacle configurations and the randomness of users.
VI Conclusion
A novel method was proposed to tackle typical practical requirements of an RIS-assisted system, such as exploiting the user distribution to maximize performance, exploiting the obstacle configuration to maximize the coverage of the system, and handling user randomness. This is the first work, to the best of our knowledge, to handle the ideal hierarchical optimization directly - maximizing coverage, while also ensuring good performance. A set of min-max solutions for the RIS placement was proposed to be obtained by searching over multiple discrete sets formed based on numerous realizations from the user distribution. Finally, a recursive strategy was used to find the final region to deploy the RIS. The superiority of the proposed algorithm was validated using simulations.